The sum of two numbers is 130. One of the numbers is 50 more than the other number. What are the two numbers?
Answer by edjones(8007)

x+y=130
x=50y
50y+y=130
51y=130
y=130/51
.
x=50 * 130/51
=6500/51
.
Ed